OpenStudy (anonymous):

Which expression is equivalent to x^-4? x^-4 • x^1 x^-2 + x^-2 (1/x)^4 x^2 – x^6

OpenStudy (marissalovescats):

Well do you know what the rule is for negative exponents?

OpenStudy (marissalovescats):



OpenStudy (marissalovescats):

So what would x^-4 be?

OpenStudy (anonymous):

(1/x)^4=1^4/x^4=1/x^4=x^(-4)

OpenStudy (anonymous):

x^-4 = 1/x^4 so the 3rd is OK